# Monotonic warpings for additive and deep Gaussian processes

### v1

@@ Abstract
Gaussian processes (GPs) are canonical as surrogates for computer experiments because they enjoy a degree of analytic tractability. But that breaks when the response surface is constrained, say to be monotonic. Here, we provide a “mono-GP” construction for a single input that is highly efficient even though the calculations are non-analytic. Key ingredients include transformation of a reference process and elliptical slice sampling. We then show how mono-GP may be deployed effectively in two ways. One is additive, extending monotonicity to more inputs; the other is as a prior on injective latent warping variables in a deep Gaussian process for (non-monotonic, multi-input) non-stationary surrogate modeling. We provide illustrative and benchmarking examples throughout, showing that our methods yield improved performance over the state-of-the-art on examples from those two classes of problems.
@@ 1 Introduction
This paper begins by taking things in the opposite direction: limiting GPs to make the most of smaller $n$ from a constrained process. However, computational tractability and scalability will remain a theme throughout, and we shall eventually turn to large-$n$ non-stationary modeling. In particular, we study GPs for modeling monotonic response surfaces. The literature here largely focuses on a single input (Tran et al.,, 2023; Ustyuzhaninov et al.,, 2020; Riihimäki and Vehtari,, 2010), motivated by case studies in materials science and health care, but has recently been extended to multiple inputs (López-Lopera et al.,, 2022).
@@
We then pivot to another application, involving injective priors over the warpings provided by the latent layers of a deep Gaussian process (DGP; Damianou and Lawrence,, 2013; Sauer et al.,, 2023c, b). DGPs are a recently popular non-stationary modeling apparatus in ML and computer experiments (Sauer et al.,, 2023a). We wish to address a DGP-related concern recently raised in spatial and geostatistical settings (Zammit-Mangion et al.,, 2022). While a more cavalier attitude – “let the posterior do whatever it wants” – can sometimes be advantageous, we agree with Zammit-Mangion et al., that non-injective warpings are less interpretable and, in some (possibly most) cases, lead to over-fitting and inferior results out-of-sample. We show that applying our mono-GP separately to the warping of each individual input of a deep GP (mw-DGP), which guarantees an injective input map by construction, is both easier to interpret than an unconstrained warping and leads to more accurate predictions for surrogate modeling of computer simulation experiments.
@@ 2.2 Elliptical slice sampling
Each thick colored line represents a subsequent $F_{n}^{(t)}$ for some $t$ Thinner, dashed lines of identical color indicate rejected samples within ESS before the thick one was accepted. Observe that it doesn’t take many iterations before one is accepted, and that the chain converges very quickly to something reasonable from a disadvantageous starting position.
@@ 3.2 Full model and posterior inference
$p(\mu,\sigma^{2})$ $\propto 1/\sigma^{2}$ improper Jeffereys prior
@@
$\hat{\mu}(\mathcal{X}_{n^{\prime}})$ $=\frac{1}{T}\sum_{t=1}^{n}\hat{\mu}^{(t)}+\nu^{(t)}(\mathcal{F}_{n^{\prime}}^{(t)}-\textstyle\frac{1}{2})$ (12)
@@ 5 Monotonic warpings for deep GPs
Deep Gaussian processes (DGPs) originate in spatial statistics (e.g., Sampson and Guttorp,, 1992; Schmidt and O’Hagan,, 2003) for non-stationary modeling (e.g., Paciorek and Schervish,, 2003; Sauer et al.,, 2023a), and have been recently re-popularized in ML (Salimbeni and Deisenroth,, 2017; Bui et al.,, 2016; Cutajar et al.,, 2017; Laradji et al.,, 2019) by analogy to deep neural networks (DNNs). Here we focus on Sauer et al., (2023c) whose fully Bayesian DGP provides the UQ necessary for many surrogate modeling applications of interest to us. Sauer et al.,’s ESS-based inferential framework laid the groundwork for our monotonic setup in Section 3.1. We provide a brief DGP review here, with the minimum detail required to explain our contribution. As a disclaimer, note that we are pivoting back to ordinary (not monotonic) regression, with mono-GP featuring as an important subroutine.
@@ 5.1 Component swap
but coded to the unit cube. A non-stationary GP is essential because there are both steep (cross) and flat (tray) regions. We train on an LHS of size $n=40$ and predict on a dense 2d grid of size $n^{\prime}=900$. Throughout our DGP-based experiments we use deepgp-package defaults (i.e., 10K MCMC samples), discarding 1K as burn-in and thinning by 10 for $T=900$ total. Basically we use fit_two_layer with and without a new option: monowarp = TRUE.
@@ 5.2 Benchmarking
Figure 13: Non-stationary benchmark out-of-sample metrics. Smaller is better for all.
@@
The second column of Figure 13 is for the Langley Glide-Back Booster (LGBB; Pamadi et al.,, 2004). The LGBB simulator models aeronautically relevant forces as a rocket booster re-enters Earth’s atmosphere as a function of its speed (mach number), angle of attack, and side slip angle. Here we study the lift force as predicted by those $p=3$ input variables. Gramacy and Lee, (2008)’s TGP was originally developed for the LGBB because it exhibits non-stationary dynamics which manifest as coordinate-wise regime changes in speed and angle of attack. Access to novel simulations for LGGB is no longer possible (previously, runs could only be obtained on NASA’s Columbia supercomputer), so here we use a corpus of 780 training runs provided by Gramacy and Lee, (2009) and subsamples of a densely interpolated set of testing $>37$K outputs provided by a TGP fit from that same paper. For more detail see Gramacy, (2020, Chapter 2). Here we use all simulation outputs as training data $(n=780)$ and a random subsample of TGP predictions for testing $(n^{\prime}=1000)$.
@@
Our final example involves simulations of atmospheric drag for the Gravity Recovery and Climate Experiment (GRACE) satellite in low-Earth orbit (Mehta et al.,, 2014) via a corpus of one million LHS-based runs provided by Sun et al., (2019). These data/simulations involve $p=7$ inputs, and here we consider randomly sub-sampled GRACE runs in pure Hydrogen of size $(n,n^{\prime})=(550,1000)$. As the model is not deterministic, fifty training runs are replicated to provide each model a measure of the uncertainty in the simulation. The results are shown in the right/third column of panels of Figure 13. In contrast to our previous examples, there is no reason to suspect that it might be advantageous to limit non-stationary modeling to axis-aligned dynamics, monotonic or otherwise, beyond the regularization that implies. In fact, observe in the figure that TGP, with its axis-aligned splits, performs worst of all in some cases. A DGP is better than both TGP and an ordinary GP, but in the latter perhaps not by an impressive margin. Nevertheless, mw-DGP is by far the best by both metrics.
@@ 6 Discussion
We have provided a new prior on monotonic Gaussian processes (GPs) for a single input variable that we illustrate is valuable for two disparate classes of problems: additive monotonic GPs for multiple inputs and as intermediate warpings for deep GPs (DGPs). There are two key ingredients that make this work well and which distinguish it from earlier approaches to mono-GP. One is a reference process that limits computational effort in the face of cubic computational bottlenecks, and which is essential for our cumulative-sum-based monotonic transformation. The other is elliptical slice sampling (ESS), a powerful but often overlooked MCMC mechanism that is ideally suited to high-dimensional multivariate normal sampling. We showed that our mono-GPs and our monowarped DGPs (mw-DGPs) outperform ordinary GP and DGP alternatives. In our Appendix we also show that our methods are faster, except perhaps on the smallest problems.
We speculate that other simple constraints on GPs (and DGP warpings), besides monotonicity, can be handled in a similar way, as long as a transformation can be used in conjunction with ESS. This could be particularly powerful in small-data contexts, e.g., with expensive computer simulations where much is known about the physics. It may additionally be possible to expand mono-GP to multiple inputs by some construct other than additively. For example, it may be worth exploring active subspace-based GPs (Wycoff,, 2023) in order to rotate into an appropriate axis-aligned regime.
@@ A.1 Choosing reference grid density
Observe in the left-most panel, corresponding to $n_{g}=10$, that the linear interpolation in the mean/black line is visibly detectable. It may also be visible in the second panel, corresponding to $n_{g}=20$, depending on screen resolution/magnification or printer quality. When we zoom in on a high-quality laptop screen we can see the line segments in the $n_{g}=20$ plot. But this is not the case for either of the the right pair of panels.

### v2

@@ Abstract
Gaussian processes (GPs) are canonical as surrogates for computer experiments because they enjoy a degree of analytic tractability. But that breaks when the response surface is constrained, say to be monotonic. Here, we provide a “mono-GP” construction for a single input that is highly efficient even though the calculations are non-analytic. Key ingredients include transformation of a reference process and elliptical slice sampling. We then show how mono-GP may be deployed effectively in two ways. One is additive, extending monotonicity to more inputs; the other is as a prior on injective latent warping variables in a deep Gaussian process for (non-monotonic, multi-input) nonstationary surrogate modeling. We provide illustrative and benchmarking examples throughout, showing that our methods yield improved performance over the state-of-the-art on examples from those two classes of problems.
@@ 1 Introduction
This paper begins by taking things in the opposite direction: limiting GPs to make the most of smaller $n$ from a constrained process. However, computational tractability and scalability will remain a theme throughout, and we shall eventually turn to large-$n$ nonstationary modeling. In particular, we study GPs for modeling monotonic response surfaces. The literature here largely focuses on a single input (Tran et al.,, 2023; Ustyuzhaninov et al.,, 2020; Riihimäki and Vehtari,, 2010; Golchi et al.,, 2015), motivated by case studies in materials science and health care, but has recently been extended to multiple inputs (López-Lopera et al.,, 2022).
@@
We then pivot to another application, involving injective priors over the warpings provided by the latent layers of a deep Gaussian process (DGP; Damianou and Lawrence,, 2013; Sauer et al.,, 2023c, b). DGPs are a recently popular nonstationary modeling apparatus in ML and computer experiments (Sauer et al.,, 2023a). We wish to address a DGP-related concern recently raised in spatial and geostatistical settings (Zammit-Mangion et al.,, 2022). While a more cavalier attitude – “let the posterior do whatever it wants” – can sometimes be advantageous, we agree with Zammit-Mangion et al., that non-injective warpings are less interpretable and, in some (possibly most) cases, lead to over-fitting and inferior results out-of-sample. We show that applying our mono-GP separately to the warping of each individual input of a deep GP (mw-DGP), which guarantees an injective input map by construction, is both easier to interpret than an unconstrained warping and leads to more accurate predictions for surrogate modeling of computer simulation experiments.
@@ 2.2 Elliptical slice sampling
Each thick colored line represents a subsequent $F_{n}^{(t)}$ for some $t$. Thinner, dashed lines of identical color indicate rejected samples within ESS before the thick one was accepted. Observe that it doesn’t take many iterations before one is accepted, and that the chain converges very quickly to something reasonable from a disadvantageous starting position.
@@ 3.2 Full model and posterior inference
$p(\mu,\sigma^{2})$ $\propto 1/\sigma^{2}$ improper Jeffreys prior
@@
$\hat{\mu}(\mathcal{X}_{n^{\prime}})$ $=\frac{1}{T}\sum_{t=1}^{T}\hat{\mu}^{(t)}+\nu^{(t)}(\mathcal{F}_{n^{\prime}}^{(t)}-\textstyle\frac{1}{2})$ (12)
@@ 4.2 Implementation and benchmarking
Figure 11: Diagonal visual (left) and metrics on the 10d Lopez–Lopera MC experiment.
@@ 5 Monotonic warpings for deep GPs
Deep Gaussian processes (DGPs) originate in spatial statistics (e.g., Sampson and Guttorp,, 1992; Schmidt and O’Hagan,, 2003) for nonstationary modeling (e.g., Paciorek and Schervish,, 2003; Sauer et al.,, 2023a), and have been recently re-popularized in ML (Salimbeni and Deisenroth,, 2017; Bui et al.,, 2016; Cutajar et al.,, 2017; Laradji et al.,, 2019) by analogy to deep neural networks (DNNs). Here we focus on Sauer et al., (2023c) whose fully Bayesian DGP provides the UQ necessary for many surrogate modeling applications of interest to us. Sauer et al.,’s ESS-based inferential framework laid the groundwork for our monotonic setup in Section 3.1. We provide a brief DGP review here, with the minimum detail required to explain our contribution. As a disclaimer, note that we are pivoting back to ordinary (not monotonic) regression, with mono-GP featuring as an important subroutine.
@@ 5.1 Component swap
but coded to the unit cube. A nonstationary GP is essential because there are both steep (cross) and flat (tray) regions. We train on an LHS of size $n=40$ and predict on a dense 2d grid of size $n^{\prime}=900$. Throughout our DGP-based experiments we use deepgp-package defaults (i.e., 10K MCMC samples), discarding 1K as burn-in and thinning by 10 for $T=900$ total. Basically we use fit_two_layer with and without a new option: monowarp = TRUE.
@@ 5.2 Benchmarking
Figure 13: Nonstationary benchmark out-of-sample metrics. Smaller is better for all.
@@
The second column of Figure 13 is for the Langley Glide-Back Booster (LGBB; Pamadi et al.,, 2004). The LGBB simulator models aeronautically relevant forces as a rocket booster re-enters Earth’s atmosphere as a function of its speed (mach number), angle of attack, and side slip angle. Here we study the lift force as predicted by those $p=3$ input variables. Gramacy and Lee, (2008)’s TGP was originally developed for the LGBB because it exhibits nonstationary dynamics which manifest as coordinate-wise regime changes in speed and angle of attack. Access to novel simulations for LGBB is no longer possible (previously, runs could only be obtained on NASA’s Columbia supercomputer), so here we use a corpus of 780 training runs provided by Gramacy and Lee, (2009) and subsamples of a densely interpolated set of testing $>37$K outputs provided by a TGP fit from that same paper. For more detail see Gramacy, (2020, Chapter 2). Here we use all simulation outputs as training data $(n=780)$ and a random subsample of TGP predictions for testing $(n^{\prime}=1000)$.
@@
Our final example involves simulations of atmospheric drag for the Gravity Recovery and Climate Experiment (GRACE) satellite in low-Earth orbit (Mehta et al.,, 2014) via a corpus of one million LHS-based runs provided by Sun et al., (2019). These data/simulations involve $p=7$ inputs, and here we consider randomly sub-sampled GRACE runs in pure Hydrogen of size $(n,n^{\prime})=(550,1000)$. As the model is not deterministic, fifty training runs are replicated to provide each model a measure of the uncertainty in the simulation. The results are shown in the right/third column of panels of Figure 13. In contrast to our previous examples, there is no reason to suspect that it might be advantageous to limit nonstationary modeling to axis-aligned dynamics, monotonic or otherwise, beyond the implicit regularization that implies. In fact, observe in the figure that TGP, with its axis-aligned splits, performs worst of all in some cases. A DGP is better than both TGP and an ordinary GP, but in the latter perhaps not by an impressive margin. Nevertheless, mw-DGP is by far the best by both metrics.
@@ 6 Discussion
We have provided a new prior on monotonic Gaussian processes (GPs) for a single input variable that we illustrate is valuable for two disparate classes of problems: additive monotonic GPs for multiple inputs and as intermediate warpings for deep GPs (DGPs). There are two key ingredients that make it work well and which distinguish it from earlier approaches to mono-GP. One is a reference process that limits computational effort in the face of cubic computational bottlenecks, and which is essential for our cumulative-sum-based monotonic transformation. The other is elliptical slice sampling (ESS), a powerful but often overlooked MCMC mechanism that is ideally suited to high-dimensional multivariate normal sampling. We showed that our mono-GPs and our monowarped DGPs (mw-DGPs) outperform ordinary GP and DGP alternatives. In our Appendix we also show that our methods are faster, except perhaps on the smallest problems.
We note that many other statistical methods exist to model monotonic response surfaces (Tutz and Leitenstorfer,, 2007; Cai and Dunson,, 2007; Bornkamp and Ickstadt,, 2009), but the literature on imposing such a constraint for GPs is more limited. Moreover, our method fits nicely within the context of other frameworks, such as DGPs. We speculate that other simple constraints on GPs (and DGP warpings), besides monotonicity, can be handled in a similar way, as long as a transformation can be used in conjunction with ESS. This could be particularly powerful in small-data contexts, e.g., with expensive computer simulations where much is known about the physics.
As currently implemented, mono-GP does not perform as well when the response surface is not coordinate-wise additive. [See Appendix B.5.] Future research could expand mono-GP to multiple inputs by some construct other than additivity. For example, it may be worth exploring active subspace-based GPs (Wycoff,, 2023) in order to rotate into an appropriate axis-aligned regime. Additionally, mono-GP could integrate previous work using Gaussian process projection (Lin and Dunson,, 2014) that generalizes beyond the class of additive models.
@@ A.1 Choosing reference grid density
Observe in the left-most panel, corresponding to $n_{g}=10$, that the linear interpolation in the mean/black line is visibly detectable. It may also be visible in the second panel, corresponding to $n_{g}=20$, depending on screen resolution/magnification or printer quality. When we zoom in on a high-quality laptop screen we can see the line segments in the $n_{g}=20$ plot. But this is not the case for either of the right pair of panels.
@@
We remark that it may be interesting to explore non-uniform reference grids of varying size, akin to knot placement in adaptive quadrature. While this could be advantageous, it would require knowledge of the response surface, presenting a chicken-or-egg problem. We have yet to encounter any scenarios where a uniform grid of modest size is insufficient. If a denser grid is warranted in some regions, it would be easiest to increase the density of the entire grid, which only marginally affects runtime.
@@ B.4 Additional sensitivity visuals
B.5 Non-additive monotonic response surfaces
We have demonstrated that our method outperforms its competitors when the response surface is constructed additively, as the title of our paper suggests. However, at the direction of a referee we explore here an expanded family of functions that adhere to monotonicity in each dimension, but are not necessarily additive. In short, we wish to explore how well mono-GP performs when the model is mis-specified. Consider the “plateau” function
$f(x)=2\cdot\Phi\left(\sqrt{2}\left(-4-3\sum_{i=1}^{d}x_{i}\right)\right)-1\quad\textrm{for}\;x\in[-2,2]^{d}\;\left(\textrm{scaled to}\;x\in[0,1]^{d}\right),$ (14)
as displayed in the left panel of Figure 22. Note that the response is monotonically increasing in both $x_{1}$ and $x_{2}$, but it’s clear from Eq. (14) that the function is not coordinate-wise additive. More succinctly, it is monotone along the $d$-dimensional diagonal. To assess performance of our method in this context, we conduct a MC experiment comparing mono-GP to an ordinary GP, ordinary DGP, and mw-DGP on 30 unique training/testing sets with $(n,n^{\prime})=(100,1600)$. Results are shown in the middle and right panels of Figure 22.
Figure 22: Plateau function (left); Performance metrics on 30 training/test sets of size $n=100$ and $n^{\prime}=1600$ (middle, right).
The ordinary DGP wins both in terms of predictive accuracy and UQ. Interestingly, mw-DGP performs similarly to the ordinary GP, indicating that the ordinary DGP relies on relationships across dimensions to appropriately warp the inputs. By a clear margin, mono-GP falls behind the other three. Our proposed method, therefore, does not appear to excel in generality. While we note that this example is as pathological as one can get for a monotonic, non-additive surface, it illustrates that mono-GP needs modification to account for non-axis-aligned changes in the response surface or some type of pre-processing step to rotate the space into an axis-aligned format. This is a subject for future work.